The width of a rectangular field is 8ft. shorter than its length. The length is represented by l.
a. Write an expression in terms of l that represents the perimeter of the rectangle.
b. Given that the length of the rectangular field is 32 ft., find the perimeter of the field. How many 4 ft. sections of fence are required to enclose the field?
c. If the perimeter of the field is 80 ft., find the area enclosed by the fence. Support your answer.
I have been doing this all day, please help!!!

Answer by nerdybill(7384)

The width of a rectangular field is 8ft. shorter than its length. The length is represented by l.
a. Write an expression in terms of l that represents the perimeter of the rectangle.
Let L = length
then
L-8 = width
.
Perimeter = 2(L + L-8) = 2(2L-8) = 4L-16
.
b. Given that the length of the rectangular field is 32 ft., find the perimeter of the field.
Perimeter = 4L-16
Perimeter = 4(32)-16
Perimeter = 128-16
Perimeter = 112 feet
How many 4 ft. sections of fence are required to enclose the field?
112/4 = 108 sections
.
c. If the perimeter of the field is 80 ft., find the area enclosed by the fence. Support your answer.
Perimeter = 4L-16
80 = 4L-16
96 = 4L
24 feet = L (length)
.
width:
L-8 = 24-8 = 16 feet
.
Area = 24*16 = 384 square feet